# Structured Learning of Compositional Sequential Interventions

### v1

@@ 3 Algorithm and Statistical Inference
This section introduces a learning algorithm, and well as ways of quantifying uncertainties in prediction. We also allow for the learning of adaptive basis functions $\phi$. As Eq. (1) does not define a generative model, which will be necessary for multiple-steps-ahead prediction, for the remainder of this section we will assume the likelihood implied by $X_{n}^{t}(d_{n}^{1:t})=f(x_{n}^{1:t-1}(d_{n}^{1:t-1}),z_{n})+\epsilon_{n}^{t}$. Here, the conditional mean $f(\cdot,\cdot)$ is given by Eq. (1) and $\epsilon_{n}^{t}\sim\mathcal{N}(0,\sigma^{2})$, an error unaffected by $D_{n}$ and other variables. Hence, $\sigma^{2}$ requires no further needs for identification results. In general, if our likelihood is defined via a finite set of estimating equations, we can parameterize it in the multilinear form analogous to (1) and repeat the analysis of the previous section. In practice, the functionals (such as the conditional expectations of $X_{t}$) used in an estimating equation are unknown. Matrix factorization methods can be applied directly by first performing a smoothing method on the data to get plug-in estimates of these functionals Squires et al. (2022), but we will adopt instead a likelihood-based approach.
@@ 3.1 Algorithm: CSI-VAE
We treat each $\beta_{n}$ as a random effects vector, giving each entry $\beta_{nl}$ an independent zero-mean Gaussian prior with variance $\sigma^{2}_{\beta}$. Along with all hyperparameters, we optimize the (marginal) log-likelihood by gradient-based optimization. With a Gaussian likelihood, the posterior and filtering distribution of each $\beta_{n}$ can be computed in closed form, but in our implementation we used a black-box amortised variational inference framework (Kingma and Welling, 2013; Mnih and Gregor, 2014; Rezende et al., 2014) anyway that can be readily put together without specialized formulas and is easily adaptable to other likelihoods. We use a mean-field Gaussian approximation with posterior mean and variances produced by a gated recurrent unit (GRU) model (Chung et al., 2014) composed with a multilayer perceptron (MLP). The approximate posterior parameters at sample with $t$ time steps are therefore $\mu_{q,\beta,n}:=\text{MLP}(\text{GRU}_{\eta_{\beta,1}}(d_{n}^{1:t},x_{n}^{1:t},z_{n})),$ and $\log\sigma_{q,\beta,n}:=\text{MLP}(\text{GRU}_{\eta_{\beta,2}}(d_{n}^{1:t},x_{n}^{1:t},z_{n}))$. Prediction for $X_{n}^{t+1:t+\Delta}$ is done by sampling $M$ trajectories, where for each trajectory we first sample a new $\beta_{n}$ from the mean-field Gaussian approximation. We set each forward $\hat{X}_{n}^{t+i}$ to be the corresponding marginal Monte Carlo average. Each basis vector $\phi_{n}^{t}$ is parameterized via another GRU model such that $\phi_{n}^{t}:=\text{MLP}(\text{GRU}_{\phi}(x_{n}^{1:t-1},z_{n}))$. Although the theory in the previous section relies on known basis functions, we could train this GRU up to an threshold time point $T_{0}$ and freeze it from that point, each $\beta_{n}$ being defined as the unit coefficient vector for this basis. In practice, we found that doing end-to-end learning provides a modest improvement, and this will be the preferred approach in the experiments. We do find that sometimes it is more stable to only use time series up to time point $T_{0}$ for the amortized $\beta$ inference. Hence, we adopt this pipeline for any results reported in this paper, unless otherwise specified. We call our method Compositional Sequential Intervention Variational Autoencoder (CSI-VAE).
@@ 3.2 Distribution-free Uncertainty Quantification
    Hold-out predictions. We have a set of historical users, $n=1,\dots,N$, whose behaviour has been observed until time $t+\Delta$. The task is to predict the behaviour at time $t+\Delta$ of a new user, $n=N+1$, who has been observed up to time $t$, i.e. to predict $X_{N+1}^{t+\Delta}$ given $X_{N+1}^{1:t}$ and $D_{N+1}^{1:t}$. If we assume we have used the history of the new user $X_{n}^{1:t}$, $n=1,\dots,N+1$, to train the model, calibration and test samples are exchangeable.
@@
    Next-intervention predictions. We have observed the behaviour of all users, $n=1,\dots,N+1$, up to time $t$ and aim to predict the effects of the next intervention, which happens at time $t+1$ for all users, i.e. $D_{n}^{t+1}\neq 0$, holding $D_{n}^{t+2:t+\Delta}=0$. The task is to predict $X_{n}^{t+\Delta}(d_{n}^{1:t+\Delta})$ for $n=1,\dots,N+1$, but in what follows we will drop the potential outcome notation to keep notation lighter, referring to observed $D$. Calibration and test are not exchangeable, because i) the joint distribution after time $t$, $P_{T}\sim(X_{n}^{t+\Delta},X_{n}^{1:t},D_{n}^{1:t})$ may be different from the one before time $t$, $P_{C}\sim(X_{n}^{t^{\prime}+\Delta},X_{n}^{1:t^{\prime}},D_{n}^{1:t^{\prime}})$, $t^{\prime}<t-\Delta$ and ii) we only used $X_{n}^{t^{\prime}},D_{n}^{t^{\prime}}$, $t^{\prime}=1,\dots t$, for training.
CP algorithms are applied on top of given point-prediction models. In our case, the underlying model is $f$, which predicts the expected user behaviour at time $t+\Delta$ given the user history up to time $t$, i.e. $\hat{X}_{n}^{t+\Delta}:=f(D_{n}^{1:t},X_{n}^{1:t},Z_{n})\approx{\mathbb{E}}[X_{n}^{t+\Delta}~{}|~{}D_{n}^{1:t},X_{n}^{1:t},Z_{n}]$.
@@
In the proof, we use the likelihood-ratio-weighting approach of Tibshirani et al. (2019) to obtain the empirical test distribution from the calibration samples and bound its quantile from below. The statement follows from standard inequalities on the mean and variance of the $\epsilon$-perturbed distribution. We prefer this approach to conformal prediction adaptive models for time series, e.g. Gibbs and Candes (2021) or Angelopoulos et al. (2024), for two reasons: i) adaptive schemes have asymptotic coverage guarantees that can not be used to estimate the uncertainty on a single time step and ii) optimised density estimates are a byproduct of our prediction model.
@@ 4 Further Related Work
Brodersen et al. (2015) leverages Bayesian structural time-series models to estimate causal effects, and motivates our exponential decay model in Eq. (3). Unlike Brodersen et al. (2015), who focus on single interventions, our model explicitly addresses the complexity arising from sequential interventions, providing a more detailed perspective on the dynamic interplay of treatments over time. Several approaches for conformal prediction in causal inference have emerged in recent years (e.g., Tibshirani et al., 2019; Lei and Candès, 2021), including matrix completion Gui et al. (2023) and synthetic controls Chernozhukov et al. (2021). Our focus has not been on individual nor average treatment effects, but directly on expected potential outcomes, similarly to the work on causal matrix completion Athey et al. (2021); Agarwal et al. (2023b, c); Squires et al. (2022). Unlike the matrix completion literature, we focused on prediction problems, out-of-sample for both units and time steps. Moreover, the matrix completion literature is usually framed in terms of marginal expectations $\mathbb{E}[X_{n}^{t}(d)]$, as opposed to conditional expectations (notice that Agarwal et al. (2023c) considers covariates, but those are included as part of the generative model). Marginal models have some advantages when modeling multiple-step-ahead effects Evans and Didelez (2024), but they also involve complex computational considerations that we leave for future work. Finally, there is a rich literature on confounding adjustment where exogeneity of $D_{n}^{t}$ cannot be assumed, see Hernán and J (2020) for a textbook treatment. Some of the causal matrix completion literature takes the perspectiven that factors such as $\beta_{n}$ could be interpreted as latent factors conditional on which ignorability could be assumed. However, we consider this assumption hard to justify due to the unknown nature of $\beta_{n}$. We can rely on standard approaches of sequential ignorability Robins (1997) to justify our method in the absence of randomization.
@@ 5 Experiments
We run a number of synthetic and semi-synthetic experiments to evaluate the performance of the CSI-VAE approach. In this section, we summarize our experimental set-up and main results. The code for reproducing all results and figures is available online [^2] The code will be released upon acceptance, and is available in the supplementary material.; in Appendix C, we provide a detailed description of the datasets and models; in Appendix D, we present further analysis and more results; and in Appendix E we present a demo for uncertainty quantification results.
@@
Model CSI-VAE-2 CSI-VAE-3 GRU-0 GRU-1 GRU-2 CSI-VAE-2 CSI-VAE-3 GRU-0 GRU-1 GRU-2
CSI-VAE-1 $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$
@@
Datasets and oracular simulators.
The first step to assess intervention predictions is to build a set of proper ground truth test beds, where we can control different levels of combinations of interventions. We build two sets of oracular simulators. (1) The Fully-synthetic simulator is constructed primarily based on random models following our parameterization in Section 2. (2) The Semi-synthetic simulator is constructed based on a real-world dataset from Spotify [^3] https://open.spotify.com/ which is aims to predict skip behaviour of users based on its past interaction history. See details in Appendix C. For each type of simulator, we generate $5$ different versions with different random seeds. From simulator (1), we construct a simulated datasets of size $50,000$, containing $5$ different interventions happening to the units at any time after an initial burn-in period of $T_{0}=10$, although only maximum $3$ of these can be observed for any given unit, with $r=5$ being the true dimensions of the model. For simulator (2), we construct simulated datasets of size $3,000$, again containing $5$ different interventions, an initial period $T_{0}=25$, a maximum 3 different interventions per unit and $r=10$. For making predictions, we predict outcomes for interventions not applied yet within any given unit (i.e. at least from the $2$ options left). In simulator (2) parameters $\phi$ and $\beta$ are learned from real-world data. Interventions are artificial, but inspired by the process of showing different proportions of track types to an user in a Spotify-like environment. For both setups, we use a data ratio of $0.7,0.1,0.2$ for training, validation and test, respectively. We report the final results in Table 1 with another constructed holdout set ($5,000$ for fully-synthetic and $1,000$ for semi-synthetic).
Compared models.
We implemented three variations: (1) CSI-VAE-1 follows exactly our setup in Section 2; (2) CSI-VAE-2 can be considered as an ablation study, which relaxes the product form of Eq. 1 and replace it with a black-box MLP applied directly to $(\phi_{n}^{t},\beta_{n},\psi_{n}^{t})$ (and hence may not guarantee identifiability); (3) CSI-VAE-3 as another ablation study, where the equation for $\psi$ (Eq. (2)) is replaced by a black-box function taking the sequence of interventions $D$ as a standard time series. We compare our model against: (1) GRU-0, a black-box gated recurrent unit (GRU, Cho et al. (2014)) composed with a MLP, using only the past history of $X_{n}^{1:t}$ and $Z_{n}$; (2) GRU-1, another GRU composed with a MLP that takes into account also the latest intervention $D_{n}^{t}$; and (3) GRU-2, which uses not only $D_{n}^{t}$ but also the entire past history $D_{n}^{1:t}$ just like CSI-VAE. In general, GRU-2 can be considered as a very strong and generic black-box baseline model.
@@ Results.
Each experiment was repeated 5 times, using Adam Kingma and Ba (2015) at a learning rate of 0.01, with 50 epochs in the fully-synthetic case and 100 for the semi-synthetic, which was enough for convergence. We selected the best iteration point based on small holdout set. The main results are presented in Tables 1 and 2, which shows the superiority of our model against the strong baselines. We also observed that the identifiable results and compositional interactions of intervention effect are both critical, as evidenced by the drop in performance for CSI-VAE 2 and 3 in Table 1. In Appendix D, we provide the following further experiments: (1) different choices of $r$ (to summarize: using $r$ less than the true value gracefully underfits, while there is evidence of small overfitting for choices of $r$ which overshoot the true value but where we can enforce regularization techniques; (2) different size of the training data (to summarize: even with more data provided, our model consistently outperforms the blackbox models by a significant margin. We show that a generic blackbox cannot solve this problem by simply feeding in more data); and (3) a demonstration of conformal prediction that allow us to better calibrate the predictive coverage compared to vanilla model-based prediction.

### v2

@@ 3 Algorithm and Statistical Inference
This section introduces a learning algorithm, as well as ways of quantifying uncertainty in prediction. We also allow for the learning of adaptive basis functions $\phi$. As Eq. (1) does not define a generative model, which will be necessary for multiple-steps-ahead prediction, for the remainder of this section we will assume the likelihood implied by $X_{n}^{t}~{}|~{}x_{n}^{1:t-1},z_{n},\mathrm{do}(d_{n}^{1:t})\sim\mathcal{N}(f(x_{n}^{1:t-1},z_{n},d_{n}^{1:t}),\sigma^{2})$. Here, the conditional mean $f(\cdot,\cdot,\cdot)$ is given by Eq. (1). As $\sigma^{2}$ is not affected by $D$, it can be easily shown to not require further identification results. In general, if parameters in our likelihood are implied by a finite set of estimating equations, we can parameterize it in the multilinear form analogous to (1) and repeat the analysis of the previous section. In practice, the functionals (such as the conditional expectations of $X_{t}$) used in an estimating equation are unknown. Matrix factorization methods can be used directly by first applying a smoothing method to the data to get plug-in estimates of these functionals Squires et al. (2022), but we will adopt instead a likelihood-based approach.
@@ 3.1 Algorithm: CSI-VAE
We treat each $\beta_{n}$ as a random effects vector, giving each entry $\beta_{nl}$ an independent zero-mean Gaussian prior with variance $\sigma^{2}_{\beta}$. Along with all hyperparameters, we optimize the (marginal) log-likelihood by gradient-based optimization. With a Gaussian likelihood, the posterior and filtering distribution of each $\beta_{n}$ can be computed in closed form, but in our implementation we used a black-box amortized variational inference framework (Kingma and Welling, 2014; Mnih and Gregor, 2014; Rezende et al., 2014) anyway that can be readily put together without specialized formulas and is easily adaptable to other likelihoods. We use a mean-field Gaussian approximation with posterior mean and variances produced by a gated recurrent unit (GRU) model (Chung et al., 2014) composed with a multilayer perceptron (MLP). Therefore, the approximate posterior at $t$ time steps follows from $\mu_{q,\beta,n}:=\text{MLP}(\text{GRU}_{\eta_{\beta,1}}(d_{n}^{1:t},x_{n}^{1:t},z_{n})),$ and $\log\sigma_{q,\beta,n}:=\text{MLP}(\text{GRU}_{\eta_{\beta,2}}(d_{n}^{1:t},x_{n}^{1:t},z_{n}))$. Prediction for $X_{n}^{t+1:t+\Delta}$ is done by sampling $M$ trajectories, where for each trajectory we first sample a new $\beta_{n}$ from the mean-field Gaussian approximation. We set each forward $\hat{X}_{n}^{t+i}$ to be the corresponding marginal Monte Carlo average. Each basis vector $\phi_{n}^{t}$ is parameterized via another GRU model such that $\phi_{n}^{t}:=\text{MLP}(\text{GRU}_{\phi}(x_{n}^{1:t-1},z_{n}))$. Although the theory in the previous section relies on known basis functions, we could train this GRU up to a threshold time point of $T_{0}$ and freeze it from that point, each $\beta_{n}$ being defined as the unit-level coefficient vector under this basis. In practice, we found that doing end-to-end learning provides a modest improvement, and this will be the preferred approach in the experiments. We do find that it is sometimes more stable to condition only up to time $T_{0}$ for the computation of the variational posterior of $\beta$. Hence, we adopt this pipeline for any results reported in this paper, unless otherwise specified. We call our method Compositional Sequential Intervention Variational Autoencoder (CSI-VAE).
@@ 3.2 Distribution-free Uncertainty Quantification
    Hold-out predictions. We have a set of historical users, $n=1,\dots,N$, whose behavior has been observed until time $t+\Delta$. The task is to predict the behaviour at time $t+\Delta$ of a new user, $n=N+1$, who has been observed up to time $t$, i.e. to predict $X_{N+1}^{t+\Delta}$ given $X_{N+1}^{1:t}$ and $D_{N+1}^{1:t}$. If we assume we have used the history of the new user $X_{n}^{1:t}$, $n=1,\dots,N+1$, to train the model, calibration and test samples are exchangeable.
@@
    Next-intervention predictions. We have observed the behavior of all users, $n=1,\dots,N+1$, up to time $t$ and aim to predict the effects of the next intervention, which happens at time $t+1$ for all users, i.e. $D_{n}^{t+1}\neq 0$, holding $D_{n}^{t+2:t+\Delta}=0$. The task is to predict $X_{n}^{t+\Delta}$ under $\mathrm{do}(d_{n}^{1:t+\Delta})$ for $n=1,\dots,N+1$, but in what follows we will drop the explicit $\mathrm{do}(\cdot)$ indexing to keep notation lighter, referring explicitly to past observed $D^{1:t}$ only (as its sampling distribution matters), and assuming sequential ignorability by assumption or randomization. Calibration and test are not exchangeable, because i) the joint distribution after time $t$, $P_{T}\sim(X_{n}^{t+\Delta},X_{n}^{1:t},D_{n}^{1:t})$ may be different from the one before time $t$, $P_{C}\sim(X_{n}^{t^{\prime}+\Delta},X_{n}^{1:t^{\prime}},D_{n}^{1:t^{\prime}})$, $t^{\prime}<t-\Delta$ and ii) we only used $X_{n}^{t^{\prime}},D_{n}^{t^{\prime}}$, $t^{\prime}=1,\dots t$, for training.
CP algorithms are applied on top of given point-prediction models, where we will use $\cal X$ to denote the cross-sectional sample space of any $X_{t}$. In our case, the underlying model is $f$, which predicts the expected user behaviour at time $t+\Delta$ given the user history up to time $t$, i.e. $\hat{X}_{n}^{t+\Delta}:=f(D_{n}^{1:t},X_{n}^{1:t},Z_{n})\approx{\mathbb{E}}[X_{n}^{t+\Delta}~{}|~{}D_{n}^{1:t},X_{n}^{1:t},Z_{n},\mathrm{do}(D^{1:t},d^{t+1},0^{t+2:t+\Delta})]$ for a chosen implicit $d^{t+1}$ and with $0^{t+2:t+\Delta}$ denoting the default action 0 being taken at $t+2,\dots,t+\Delta$.
@@
In the proof, we use the likelihood-ratio-weighting approach of Tibshirani et al. (2019) to obtain the empirical test distribution from the calibration samples and bound its quantile from below. The statement follows from standard inequalities on the mean and variance of the $\epsilon$-perturbed distribution. We prefer this approach to conformal prediction adaptive models for time series, e.g. Gibbs and Candes (2021) or Anastasios et al. (2024), for two reasons: i) adaptive schemes have asymptotic coverage guarantees that can not be used to estimate the uncertainty on a single time step and ii) optimized density estimates are a byproduct of our prediction model.
@@ 4 Further Related Work
Brodersen et al. (2015) leverages Bayesian structural time-series models to estimate causal effects, and motivates our exponential decay model in Eq. (3). Unlike Brodersen et al. (2015), who focus on single interventions, our model explicitly addresses the complexity arising from sequential interventions, taking a more detailed perspective on the dynamic interplay of treatments over time. Several approaches for conformal prediction in causal inference have emerged in recent years (e.g., Tibshirani et al., 2019; Lei and Candès, 2021), including matrix completion Gui et al. (2023) and synthetic controls Chernozhukov et al. (2021). Our focus has not been on individual nor average treatment effects, but directly on expected potential outcomes, similarly to the work on causal matrix completion Athey et al. (2021); Agarwal et al. (2023b, c); Squires et al. (2022). Unlike the matrix completion literature, we focused on prediction problems, out-of-sample for both units and time steps. Moreover, the matrix completion literature is usually framed in terms of marginal expectations $\mathbb{E}[X_{n}^{t}(d)]$, as opposed to conditional expectations (notice that Agarwal et al. (2023c) considers covariates, but those are included as part of the generative model). Marginal models have some advantages when modeling multiple-step-ahead effects Evans and Didelez (2024), but they also involve complex computational considerations that we leave for future work. Compositionality based on additivity instead of the factorization trick is discussed by e.g. Lotfollahi et al. (2023), but additivity lacks the connection to an explicit principle for universal function approximation. Finally, there is a rich literature on confounding adjustment where exogeneity of $D_{n}^{t}$ cannot be assumed, see Hernán and J (2020) for a textbook treatment. We can rely on standard approaches of sequential ignorability Robins (1997) to justify our method in the absence of randomization.
@@ 5 Experiments
We run a number of synthetic and semi-synthetic experiments to evaluate the performance of the CSI-VAE approach. In this section, we summarize our experimental set-up and main results. The code for reproducing all results and figures is available online [^2] The code is available at https://github.com/jialin-yu/CSI-VAE. In Appendix C, we provide a detailed description of the datasets and models. In Appendix D, we present further analysis and more results. Finally, in Appendix E, we present an illustration of uncertainty quantification results.
Datasets and oracular simulators.
The first step to assess intervention predictions is to build a set of proper ground truth test beds, where we can control different levels of combinations of interventions. We build two sets of oracular simulators. (1) The Fully-synthetic simulator is constructed primarily based on random models following our parameterization in Section 2. (2) The Semi-synthetic simulator is constructed based on a real-world dataset from Spotify [^3] https://open.spotify.com/ which aims to predict skip behavior of users based on their past interaction history. See details in Appendix C. For each type of simulator, we generate $5$ different versions with different random seeds. From simulator (1), we construct a simulated dataset of size $50,000$, containing $5$ different interventions happening to the units at any time after an initial burn-in period of $T_{0}=10$, although only maximum $3$ of these can be observed for any given unit. We set $r=5$ as the true dimensionality of the model. For simulator (2), we construct simulated datasets of size $3,000$, again containing $5$ different interventions, an initial period $T_{0}=25$, a maximum of 3 different interventions per unit, and $r=10$. The task is to predict outcomes for interventions not applied yet within any given unit (i.e., at least from the $2$ options left). In simulator (2), parameters $\phi$ and $\beta$ are learned from real-world data. Interventions are artificial, but inspired by the process of showing different proportions of track types to an user in a Spotify-like environment. For both setups, we use a data ratio of $0.7,0.1,0.2$ for training, validation and test, respectively. We report the final results in Table 1 with another constructed holdout set ($5,000$ points for the fully-synthetic case, and $1,000$ for the semi-synthetic one).
@@
Compared models.
We implemented three variations: (1) CSI-VAE-1 follows exactly our setup in Section 2; (2) CSI-VAE-2 can be considered as an ablation study, which relaxes the product form of Eq. (1) and replace it with a black-box MLP applied directly to $(\phi_{n}^{t},\beta_{n},\psi_{n}^{t})$ (and hence may not guarantee identifiability); (3) CSI-VAE-3 is another ablation study, where the equation for $\psi$ (Eq. (2)) is replaced by a black-box function taking the sequence of actions $D$ as a standard time-series. We compare our model against: (1) GRU-0, a black-box gated recurrent unit (GRU, Cho et al. (2014)) composed with a MLP, using only the past history of $X_{n}^{1:t}$ and $Z_{n}$; (2) GRU-1, another GRU composed with a MLP that takes into account also the latest intervention $D_{n}^{t}$; and (3) GRU-2, which uses not only $D_{n}^{t}$ but also the entire past history $D_{n}^{1:t}$ just like CSI-VAE. In general, GRU-2 can be considered as a very strong and generic black-box baseline model. In addition, we conduct experiments comparing other popular black-box baseline models: (4) LSTM, Hochreiter and Schmidhuber (1997); and (5) Transformer, Vaswani et al. (2017). For those, we use the same input setup as in the case of the GRU-2 model.
@@
Model CSI-VAE-2 CSI-VAE-3 GRU-0 GRU-1 GRU-2 LSTM Transformer CSI-VAE-2 CSI-VAE-3 GRU-0 GRU-1 GRU-2 LSTM Transformer
CSI-VAE-1 $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$ $<0.001^{**}$
@@
[…]
@@ Results.
Each experiment was repeated 5 times, using Adam Kingma and Ba (2015) at a learning rate of 0.01, with 50 epochs in the fully-synthetic case and 100 for the semi-synthetic, which was enough for convergence. We selected the best iteration point based on a small holdout set. The main results are presented in Tables 1 and 2, which show the superiority of our model against strong baselines. We also observed that the identifiability results and compositional interactions of intervention effects are both critical, as evidenced by the drop in performance for CSI-VAE 2 and 3 in Table 1.
In Appendix D, we provide the following further experiments: (1) different choices of $r$ (summary: using $r$ less than the true value gracefully underfits, while there is evidence of some overfitting for choices of $r$ which overshoot the true value — mitigated by regularization); (2) different sizes of the training data (summary: even with more data provided, our model consistently outperforms the black-box models by a significant margin; we show that a generic black-box cannot solve this problem by simply feeding in more data); and (3) a demonstration of conformal prediction that allow us to better calibrate the predictive coverage compared to vanilla model-based prediction.